OpenStudy (anonymous):

@IMStuck It's four questions:

OpenStudy (anonymous):

Find the derivative of f(x) = -8/x at x = 11. I got \[\frac{ 8 }{ 11 }\]

OpenStudy (imstuck):

Ok give me a sec on that one. Give me all of them, why don't you?

OpenStudy (anonymous):

Okay. Use graphs and tables to find the limit and identify any vertical asymptotes of the function. I got 1/-16 as the limit and I don't know about the vertical asymptotes.

OpenStudy (anonymous):

Find the fifth roots of 243(cos 240° + i sin 240°). This one I actually thought I had the hang of but I confused myself and I actually am completely lost.

OpenStudy (anonymous):

The last one is Find an equation in standard form for the hyperbola with vertices at (0, ±6) and asymptotes at y = ±3/7x. And I got \[\frac{ y^2 }{ 36 }-\frac{ x^2 }{ 196 }=1\]

OpenStudy (imstuck):

That first one, you're in calculus?

OpenStudy (anonymous):

I'm in pre-calc. I just remember the power rule that my teacher told me about but none of my choices have what I come up with so I thought it would be 8/11.

OpenStudy (imstuck):

From what I remember about this, you can rewrite that as \[f(x)=-8x ^{-1}\]Then take the derviative as you normally would. The exponent comes down in the front adn gets multiplied by the rest of it. Then of course you have to subtract the exponent from the one that is there already. Oh boy, how confusing! Let's look at it